### v2

# HybridVLA: Collaborative Diffusion and Autoregression in a Unified Vision-Language-Action Model

Abstract
Recent advancements in vision-language models (VLMs) for common-sense reasoning have led to the development of vision-language-action (VLA) models, enabling robots to perform generalized manipulation. Although existing autoregressive VLA methods leverage large-scale pretrained knowledge, they disrupt the continuity of actions. Meanwhile, some VLA methods incorporate an additional diffusion head to predict continuous actions, relying solely on VLM-extracted features, which limits their reasoning capabilities. In this paper, we introduce HybridVLA, a unified framework that seamlessly integrates the strengths of both autoregressive and diffusion policies within a single large language model, rather than simply connecting them. To bridge the generation gap, a collaborative training recipe is proposed that injects the diffusion modeling directly into the next-token prediction. With this recipe, we find that these two forms of action prediction not only reinforce each other but also exhibit varying performance across different tasks. Therefore, we design a collaborative action ensemble mechanism that adaptively fuses these two predictions, leading to more robust control. In experiments, HybridVLA outperforms previous state-of-the-art VLA methods across various simulation and real-world tasks, including both single-arm and dual-arm robots, while demonstrating stable manipulation in previously unseen configurations.
Figure 1: HybridVLA innovatively integrates diffusion and autoregressive action prediction within a single LLM, fully leveraging the continuity and probabilistic nature of diffusion alongside the reasoning capabilities of autoregressive modeling. It undergoes pretraining on large, diverse, cross-embodied real-world robotic datasets and is further fine-tuned on both simulation and self-collected real-world data. HybridVLA achieves remarkable performance across various tasks, demonstrating strong generalization to unseen manipulated objects, backgrounds, spatial positions, and lighting conditions.
1 Introduction
Recently, vision-language models (VLMs) [58, 3, 51, 112, 5, 21, 114, 113] have demonstrated exceptional abilities in instruction-following and common-sense reasoning, driven by pretraining on internet-scale image-text pairs. Building on this success, several studies have extended VLMs into vision-language-action (VLA) models, enabling them to generate action plans[1, 18, 34, 7] or predict SE(3) poses [10, 47, 56]. VLA models enable robots to interpret visual observations and language conditions, generating generalizable actions for control. Therefore, effectively harnessing the inherent capabilities of VLMs to develop VLA models for stable manipulation in dynamic environments is essential [55].
On the one hand, some existing VLA methods [10, 56, 47, 74] quantize continuous actions into discrete bins, replacing part of the vocabulary in large language models (LLMs). These autoregressive approaches mimic the next-token prediction of VLMs, effectively harnessing their large-scale pretrained knowledge while preserving reasoning capabilities. Although such methods enable generalized manipulation skills [47], the quantization process disrupts the continuity of action poses [97]. On the other hand, some VLA approaches [60, 33, 54, 100] incorporate a policy head (e.g., MLP or LSTM [25]) to transform LLM output embeddings into continuous action poses. However, these regressive methods overlook the scalability of the policy head and fail to incorporate probabilistic action representations [52, 14].
Building on the success of diffusion models in content generation [31, 32, 73, 80, 28], diffusion policies have recently been introduced in robotic imitation learning [14, 110, 45, 71, 79, 104]. Unlike regressive deterministic policy heads, $\pi_{0}$[8], CogACT [52], and DiVLA [97] incorporate a diffusion head after VLMs, leveraging probabilistic noise-denoising for action prediction. While diffusion-based VLA methods enable precise manipulation, the diffusion head operates independently of the VLM, relying solely on VLM-extracted features as input conditions. Consequently, it fails to fully leverage the VLM’s reasoning capabilities. Given these advantages and limitations, a question arises: “How can we elegantly construct a unified VLA model that seamlessly integrates the strengths of both autoregressive and diffusion policies, rather than simply concatenating them?
To achieve this, we propose HybridVLA, a unified framework that equips VLMs with both diffusion and autoregressive action prediction capabilities, enabling robust execution across diverse and complex manipulation tasks. Unlike previous diffusion-based VLA methods [8, 52], which append independent diffusion head after the LLM, HybridVLA seamlessly integrates diffusion modeling into the autoregressive next-token prediction within a single LLM, as shown in Figure 1. Specifically, we introduce a collaborative training recipe that encodes the diffusion-noised action as a continuous vector and projects it into the LLM’s word embedding space. To ensure consistency when combining the two generation methods, a token sequence formulation is designed to systematically organize multimodal input, diffusion action, and autoregressive action tokens while linking them through specialized marker tokens. With this design, we observe that the two action predictions not only reinforce each other but also exhibit varying performance across different tasks. For instance, the diffusion prediction excels in intricate tasks, while the autoregressive prediction performs better in tasks requiring rich semantic understanding. Therefore, a collaborative action ensemble mechanism is proposed, where the two predictions are adaptively fused based on autoregressive action token confidence, improving robustness in manipulation.
To enhance HybridVLA’s generalization, we employ a step-by-step training approach [8], first performing large-scale pretraining, followed by fine-tuning on downstream tasks, as shown at the top of Figure 1. In addition to being initialized with parameters from a pretrained VLM [44], our model undergoes further pretraining on large, diverse, cross-embodiment robotic datasets, including Open X-Embodiment [69], DROID [46], and ROBOMIND [101], encompassing 760K trajectories and over 10K A800 GPU training hours. Subsequently, HybridVLA is fine-tuned on high-quality simulation data [36] and self-collected real-world data, achieving state-of-the-art (SOTA) manipulation performance across a variety of tasks with both single-arm and dual-arm robots. Meanwhile, HybridVLA demonstrates sufficient generalization capabilities to unseen manipulated objects, backgrounds, spatial positions, and lighting conditions during real-world testing, highlighting the effectiveness of our collaborative model design and training recipe. To optimize inference speed, we introduce HybridVLA-dif, which integrates diffusion and autoregressive generation during training but relies exclusively on diffusion-based actions for inference at 9.4 Hz. In summary, our contributions are as follows:
  • We propose HybridVLA, a unified model that seamlessly integrates diffusion and autoregressive action prediction within a single LLM, effectively combining the continuous nature of diffusion-based actions with the reasoning capabilities of autoregressive generation.
  • We introduce a collaborative training recipe that bridges the gap between the two action generation approaches, enabling mutual reinforcement. Additionally, we propose a collaborative action ensemble mechanism that adaptively fuses diffusion- and autoregressive-based actions, enhancing manipulation robustness.
  • Our proposed method achieves SOTA performance across diverse tasks while demonstrating strong generalization to unseen configurations.
2 Related Work
Traditional robotic manipulation primarily relies on state-based reinforcement learning [4, 23, 42, 109], whereas recent approaches [92, 20, 9, 14] integrate visual observations for imitation learning. Building on the strong reasoning capabilities of vision-language models (VLMs) [58, 3, 51, 41, 27, 50], recent research has integrated them into robotic manipulation [53, 106, 107, 59].
Vision-language-action (VLA) models. Some studies [1, 18, 34, 35] enable robots to interpret both language and visual observations, automatically generating task plans. Meanwhile, vision-language-action (VLA) models leverage the inherent reasoning abilities of VLMs to predict low-level SE(3) poses. Specifically, RT2[10] quantizes 7-DoF actions into discrete bins for autoregressive pose prediction. Building on this, ManipLLM[56] incorporates affordance priors through chain-of-thought reasoning, while OpenVLA[47] performs large-scale pretraining on the Open X-Embodiment dataset[69], enhancing generalization capabilities. FAST [74] applies the discrete cosine transform to enable fast and scalable training of autoregressive-based VLA models. To support continuous action prediction, some VLA approaches [60, 33, 54, 100] incorporate a policy head, such as an MLP or LSTM [25], and use regression loss for imitation learning. However, quantization in autoregressive methods disrupts action continuity, while regressive methods fail to incorporate probabilistic action representations.
Diffusion models in robotics. Building on the success of diffusion models in content generation [31, 32, 73, 80], diffusion policies have been applied in robotics, including reinforcement learning [2, 96], imitation learning [14, 72, 75, 79, 104], grasping [88, 94, 102], and motion planning [38, 82]. Following this, 3D Diffusion Actor [45] and DP3[14] employ diffusion models to interpret point cloud data. Octo[91] and RDT-1B[61] enhance a transformer model with a diffusion head to predict flexible actions.
Diffusion-based VLA models. To integrate diffusion with VLMs, $\pi_{0}$[8] adds a diffusion expert head that generates actions through flow matching, while TinyVLA [98] incorporates a simple diffusion head after the lightweight VLM. CogACT[52] and DiVLA [97] decouple reasoning and action prediction into the VLM and an injected diffusion head, respectively. However, in these methods, the diffusion head functions as a separate module, relying solely on LLM-extracted features for action conditions, limiting its use of the VLM’s pretrained knowledge and reasoning capabilities. Beyond robotics, in general multimodal scenarios, some works [22, 99, 103, 105] jointly address multimodal understanding and generation, while others [115, 12, 116] incorporate diffusion into an autoregressive-based transformer. Unlike previous approaches that combine image and language generation while preserving their respective quality, HybridVLA introduces a robotic-specific collaborative training recipe that seamlessly incorporates diffusion-based action generation into autoregressive action generation within a single LLM, effectively leveraging the strengths of both methods to mutually enhance each other.
3 HybridVLA Method
Overview. Existing diffusion-based VLA methods [8, 97, 52] append a separate diffusion head after the VLM, using VLM-extracted features as conditions for the diffusion process. However, these methods fail to fully exploit the VLM’s inherent reasoning capabilities derived from internet-scale pretraining. In contrast, HybridVLA equips a single LLM with both diffusion and autoregressive action generation capabilities. To construct HybridVLA, we first describe the model architecture in Section 3.1. Since simply merging the two generation methods could cause inconsistency, we introduce a collaborative training recipe in Section 3.2. To further enhance robustness, we propose a collaborative action ensemble mechanism in Section 3.3.
Problem Statement. At time t, each demonstration consists of image observations $o_{t}$, language description $l_{t}$, and the current robot state $r_{t}$. Our model $\pi$ aims to predict action $a$ to control the robot arms, which can be formulated as:
$\pi:(o_{t},l_{t},r_{t})\rightarrow a_{t+1}$
Following [47, 52], the action $a$ represents the end-effector pose, which uses 7-DOF and 14-DOF for single-arm and dual-arm control, respectively. Each 7-DOF action includes 3-DOF for relative translation offsets ($[\Delta x,\Delta y,\Delta z]\in\mathbb{R}^{3}$), 3-DOF for rotation (Euler angles $\in\mathbb{R}^{3}$), and 1-DOF for the gripper state (open/closed $\in\mathbb{R}^{1}$). The ground truth (GT) and the model-predicted action are in SE(3), formulated as:
$a=[\Delta x,\Delta y,\Delta z,Roll,Pitch,Yaw,0/1]$
Figure 2: HybridVLA Framework. The input data, regardless of modality, is encoded and concatenated into our formatted token sequence. To integrate diffusion into the LLM, HybridVLA simultaneously projects the denoising timestep and noisy actions into the token sequence. The marker tokens, $<$BOD$>$ (beginning of diffusion) and $<$EOD$>$ (end of diffusion), are designed to bridge the two generation methods. By employing collaborative training to explicitly incorporate knowledge from both generation methods, these two action types reinforce each other and are adaptively ensembled to control the robot arms. For HybridVLA’s output, continuous actions are generated through iterative denoising, while discrete actions are produced autoregressively, all within the next-token prediction process.
3.1 HybridVLA Architecture
This section presents the architecture and workflow of HybridVLA, available in two model sizes, using 2.7B and 7B large language models (LLMs). Following [47], HybridVLA inherits a base architecture adapted from the Prismatic VLMs [44], leveraging its internet-scale pretrained parameters. We first introduce the two basic components—vision encoders and the LLM—as shown in Figure 2.
Vision encoders. HybridVLA leverages powerful vision encoder combinations, such as DINOv2 [70] and SigLIP [111], to capture rich semantic features $f_{d}\in\mathbb{R}^{B\times N_{v}\times 1024}$ and $f_{s}\in\mathbb{R}^{B\times N_{v}\times 1152}$. $B$ and $N$ represent batch size and token dimension, respectively. These features are concatenated along the channel dimension to form $f_{v}\in\mathbb{R}^{B\times N_{v}\times 2176}$, which is subsequently projected into the LLM’s word embedding via a projection layer. HybridVLA(2.7B) uses only the CLIP [77] model as its vision encoder. When processing multi-view images, a shared vision encoder extracts features, which are then concatenated along the token dimension.
LLM. HybridVLA adopts 7B LLAMA-2 [93] as LLM, responsible for multi-modal understanding and reasoning. Language prompts are encoded into embedding space $f_{l}\in\mathbb{R}^{B\times N_{l}\times 4096}$ using the pre-trained tokenizer, then concatenated with visual tokens and input into LLM. The other specially designed LLM inputs are presented in the next section, and the output tokens are processed in two ways. First, diffusion-based action ($a^{d}_{t+1}$) generation through a denoising process, where an MLP maps the tokens into the action space. Second, autoregressive-based action ($a^{ar}_{t+1}$) generation is performed through a detokenizer [47], which also computes the mean confidence ($c^{ar}_{t+1}$) of the action tokens, serving as a guiding factor for the collaborative action ensemble. For HybridVLA (2.7B), the workflow remains the same as that of HybridVLA (7B) but utilizes the 2.7B Phi-2 [39] as the LLM. In the next section, we introduce how to simultaneously equip a single LLM with diffusion and autoregressive action generation capabilities.
3.2 Collaborative Training Recipe
Directly combining diffusion and autoregressive pose prediction within a single LLM presents challenges such as instability and inconsistencies in next-token prediction. Therefore, we propose a collaborative training recipe that encompasses token sequence formulation design, hybrid objectives, and structured training stages.
Token sequence formulation design. As shown in the upper part of Figure 2, the input token sequence during training comprises not only visual and language tokens but also robot state, diffusion noise, and autoregressive action tokens. For the robot state, we integrate it into the LLM to enhance temporal consistency in action generation. Instead of discretizing the robot state and merging it with the language query [56] (Type 3 of Table 1), we employ a learnable MLP to map the robot state directly into the word embedding space, $f_{r}\in\mathbb{R}^{B\times 1\times 4096}$. The motivation is that diffusion action tokens are generated during the subsequent next-token prediction process, using all preceding tokens as conditions. Introducing discrete robot states could negatively impact the diffusion prediction of continuous actions. For diffusion actions, we predict them through a diffusion denoising process, conditioned on the previous tokens. The denoising step $i$ and noisy actions $a^{i}_{t}$ are projected into the LLM’s word embeddings through an MLP, represented as continuous vectors. To seamlessly connect previous multi-modal tokens, diffusion tokens, and subsequent discrete tokens within a single sequence, we introduce special beginning-of-diffusion ($<$BOD$>$) and end-of-diffusion ($<$EOD$>$) tokens to encapsulate the diffusion tokens. This design not only clarifies the boundaries between the diffusion and autoregressive generation but also prevents confusion in the next-token prediction process, such as diffusion tokens directly predicting masked discrete tokens (Type 2 of Table 1). For autoregressive actions, we quantize the end-effector pose into discrete bins and replace part of the vocabulary in the LLM [47], which is then tokenized into a sequence of discrete tokens. Owing to the inherent nature of next-token prediction, both the question and the answer (discrete action GT) are fed into the LLM during training, while only the question is provided during inference. If the autoregressive prediction is placed before the diffusion tokens, it would lead to action GT leakage, serving as a condition for diffusion training (Type 4 of Table 1). Consequently, we place the diffusion token at the front explicitly providing continuous knowledge for the subsequent autoregressive generation.
Type1(Ours) Type2 Type3 Type4
0.66 0.56 0.61 0.57
Table 1: Exploring and validating our proposed token sequence formulation. The model is trained using both diffusion and autoregressive generation but tested only on diffusion-based actions (HybridVLA-dif) across 10 simulation tasks.
Hybrid objectives. To simultaneously train diffusion and autoregressive action generation, we require two distinct loss functions. For diffusion part, following previous diffusion policies [14], we minimize the mean squared error between the predicted noise ($\epsilon_{\pi}$) from the VLA model and the GT noise ($\epsilon$). The loss function is defined as follows:
$L_{dif}=E_{a,i,c}||\epsilon-\epsilon_{\pi}(a^{i}_{t},i,c)||^{2}$
Where $\epsilon\sim N(0,1)$ and $c$ denotes the condition. Additionally, classifier-free guidance [30] is not used in order to ensure stable robot arm behavior [61]. For the autoregressive part, the cross-entropy loss ($L_{ce}$) is adopted to supervise the discrete output. With our designed token sequence formulation, the two losses can be seamlessly combined for collaborative penalization, defined as:
$L_{hybrid}=L_{dif}+L_{ce}$
Structured training stage. After loading the pretrained VLM parameters, HybridVLA undergoes two training stages with hybrid objectives: large-scale pretraining on open-source robotic data and fine-tuning on self-collected simulation and real-world data. During pretraining, we train HybridVLA for 5 epochs on 35 datasets from Open X-Embodiment [69], DROID [46], ROBOMIND [101], and others. The pretrain datasets contain 760k robot trajectories, comprising 33m frames. Due to dataset differences, pretraining relies solely on single 2D observations, whereas fine-tuning relies on either single or multi-view observations, depending on the downstream task. The details of the pretraining dataset are provided in Appendix A.1.
3.3 Collaborative Action Ensemble
During inference, given visual, language, and robot state inputs, HybridVLA concurrently generates actions using both diffusion and autoregressive methods, and then ensembles the actions to produce more stable execution.
Autoregressive actions. As shown in Figure 2, the autoregressive generation begins after the special token $<$EOD$>$. Similar to [47, 56], the generation of 7-DoF or 14-DoF actions closely mirrors the text generation process in an LLM. Unlike previous autoregressive VLA methods, HybridVLA’s autoregressive generation additionally conditions on the inherently continuous nature of actions from diffusion tokens, outperforming independent autoregressive approaches, as demonstrated in the ablation study.
Diffusion actions. When generating diffusion actions, we append the special token $<$BOD$>$ after the previous condition tokens to indicate that the model should perform the denoising process. We employ DDIM [89] sampling with $n$ sampling steps. In HybridVLA, we find that integrating diffusion into the next-token prediction process not only improves action precision by fully leveraging the reasoning abilities and pretrained knowledge of VLMs but also mitigates performance degradation when reducing the number of inference denoising steps (e.g., $n=4$), as demonstrated in the ablation study. To accelerate the sampling process, we introduce the KV cache before the diffusion tokens, forwarding conditional information, the denoising timestep, and pure noise only during the initial sampling step. In subsequent steps, the cached keys and values from the first pass are reused, while only the timestep and noise are iteratively forwarded. This strategy eliminates redundant computations and improves inference speed.
Close Close Toilet Sweep Close Phone Umbrella Frame Wine at Water Mean Infer.
Models box laptop lid seat down to dustpan fridge on base out off hanger rack plants S.R.$\uparrow$ speed
ManipLLM (7B) [56] 0.50 0.80 0.40 0.20 0.80 0.35 0.10 0.25 0.15 0.20 0.38 2.2 Hz
OpenVLA (7B) [47] 0.65 0.40 0.75 0.60 0.80 0.20 0.35 0.15 0.10 0.10 0.41 6.3 Hz
$\pi_{0}$ (2.6B) [8] 0.90 0.60 1.00 0.30 0.90 0.25 0.35 0.75 0.05 0.45 0.55 13.8 Hz
CogACT (7B) [52] 0.80 0.85 0.90 0.65 0.90 0.50 0.60 0.35 0.25 0.25 0.60 9.8 Hz
HybridVLA-dif (7B) 0.85 0.75 1.00 0.80 0.95 0.50 0.50 0.30 0.70 0.25 0.66 9.4 Hz
HybridVLA (7B) 0.85 0.95 1.00 0.90 1.00 0.50 0.50 0.70 0.50 0.50 0.74 6.1 Hz
HybridVLA (2.7B) 1.00 0.80 0.90 0.80 0.90 0.25 0.20 0.45 0.25 0.25 0.58 12.3 Hz
Table 2: Comparison of our proposed method and baselines on RLBench. We train all methods in the Multi-task setting [87] and report the success rates (S.R.) for each task. The success condition follows the definition in RLBench. “HybridVLA-dif” refers to action prediction relying solely on the diffusion process. (7B), (2.7B), and (2.6B) refer to the sizes of the LLM used in the VLA model.
Ensembled actions. After obtaining the two types of actions under our collaborative training recipe, we empirically observe two phenomena. 1) Different action types demonstrate varying performance across tasks. Diffusion-based predictions excel in precise manipulation tasks, such as Phone on base and Close laptop lid, while autoregressive predictions perform better in tasks requiring scene semantic reasoning, such as Water plants and Frame off hanger. 2) The confidence of autoregressive action tokens serves as a reliable indicator of action quality. In over 80% of successfully completed test samples, the average confidence of autoregressive action tokens exceeds 0.96. Quantitative evaluations are provided in Appendix B.1 and B.2. Therefore, as shown at the end of Figure 2, we use the mean confidence of autoregressive tokens ($c^{ar}_{t+1}$) to guide the action ensemble. If the confidence exceeds $\theta$ ($\theta=0.96$), we consider the autoregressive action ($a^{ar}_{t+1}$) sufficiently accurate and perform an average operation with the diffusion action ($a^{d}_{t+1}$). Otherwise, we rely solely on the diffusion action to control the robot. Additionally, to accelerate inference, HybridVLA-dif relies solely on diffusion-based action generation while still collaboratively learning both action generation types during training to enhance mutual reinforcement.
4 Experiment
In Section 4.1, we compare the manipulation ability and inference speed of HybridVLA with previous VLA methods in simulation environments. The effectiveness of each component is validated in Section 4.2 and Appendix B. In Section 4.3, we present both quantitative and qualitative manipulation results of HybridVLA in real-world scenarios, including single-arm and dual-arm robot tasks. The generalization capabilities of HybridVLA are examined in Section 4.4, testing on unseen manipulated instances, background, spatial positions, and lighting conditions.
4.1 Simulation Experiment
Simulation benchmark. To systematically evaluate, we select the RLBench [36] benchmark in the CoppeliaSim simulator, which contains 10 different tabletop tasks. These tasks, performed using a Franka Panda robot and a front-view camera, include Close box, Close Laptop, Toilet seat down, Sweep to dustpan, Close fridge, Phone on base, Take umbrella out, Frame off hanger, Wine at rack, and Water plants. The data are collected using pre-defined waypoints and the Open Motion Planning Library [90]. Following the frame-sampling method used in previous works [87, 24, 40], we construct the training dataset, with each task consisting of 100 trajectories.
Training and Evaluation Details. We compare our method against four previous SOTA VLA models, including autoregressive-based approaches such as ManipLLM [56] and OpenVLA [47], as well as diffusion-based methods like $\pi_{0}$[8] and CogAct[52] with a DiT-base action head. Meanwhile, we categorize our method into three modes: HybridVLA (7B), HybridVLA (2.7B), and HybridVLA-dif (7B). All modes are jointly trained using our proposed collaborative training recipe; however, HybridVLA-dif relies solely on diffusion-based action generation during inference. To ensure a fair comparison, we load the official pretrained parameters provided by each method, adhering to their respective training settings. For HybridVLA, the single-view RGB input is resized to $224\times 224$, and the robot state is consistent with predicted actions (7-DOF end-effector poses). During training, we use the AdamW optimizer with a constant learning rate of 2e-5. As shown in Figure 2, we only update the LLM and injected MLP parameters. Our models are trained for 300 epochs on 8 NVIDIA A800 GPUs with mixed-precision training. For evaluation, following [47, 52], all methods are tested with 20 rollouts from the latest epoch.
Quantitative Results. As shown in Table 2, HybridVLA(7B) achieves an average success rate of 74% across 10 distinct tasks, outperforming the previous SOTA autoregressive-based VLA (OpenVLA) and diffusion-based VLA (CogACT) by 33% and 14%, respectively. These results demonstrate that our method seamlessly introduces diffusion modeling within the LLM and effectively combines the advantages of two types of pose prediction. Remarkably, compared to CogACT and $\pi_{0}$, HybridVLA-dif also achieves performance improvements of 6% and 11%, respectively. These results highlight that, unlike previous approaches that attach the diffusion head after the VLM, our method effectively leverages the inherent capabilities of the VLM to unlock the full potential of diffusion-based action generation within the VLA model. Finally, HybridVLA(2.7B) delivers satisfactory results, confirming our method’s effectiveness in enhancing VLM manipulation capabilities across different model sizes.
Inference Speed. By harnessing VLM’s reasoning capabilities and conducting large-scale robotic pretraining, our method reduces denoising steps to 4 without any performance degradation, as shown in Figure 3. In Table 2, when tested on an NVIDIA 4090D GPU, HybridVLA-dif (7B) and HybridVLA (2.7B) achieve satisfactory frequencies comparable to CogACT (7B) and $\pi_{0}$ (2.6B), thanks to the reduced denoising steps and efficient KV cache of our method. All models are loaded with bfloat16 precision for inference. Notably, while KV cache has been used in previous autoregressive VLA methods [47, 56], we first integrate it into the LLM’s diffusion-based action prediction.
[…]
Ex0 0.74
Ex1 - - 0.66
Ex2 - - - 0.60
Ex3 - - 0.62
Ex4 - - - 0.57
Ex5 - 0.22
Ex6 - - 0.10
Ex7 - - - 0.57
Table 3: Impact of each component. AR and Dif represent autoregressive and diffusion-based action generation, respectively, using LLM as backbone. $L_{Hybrid}$ means jointly training the model by Hybrid objectives. LSP denotes large-scale pretraining on assembled robotic datasets, while RSE refers to our injected robot state embedding. CTR and CAE mean our proposed collaborative training recipe and collaborative action ensemble method.
4.2 Ablation Study
We perform ablation experiments on the 10 RLBench tasks and compute the average manipulation accuracy. The impact of each component. As shown in Table 3, a comparison between Ex1 and Ex2, as well as Ex3 and Ex4, demonstrates that under our proposed collaborative training of diffusion and autoregressive action generation, HybridVLA-dif (Ex1) and HybridVLA-ar (Ex2) outperform their individually trained counterparts. These results further confirm that our method effectively integrates diffusion’s continuous prediction capability with the reasoning ability of the autoregressive approach, enabling mutual reinforcement. Compared to Ex0, Ex5 and Ex6 highlight the essential role of large-scale pretraining and robot state injection in ensuring stable prediction. A comparison of Ex7 and Ex1 shows that simply combining diffusion and autoregressive approaches in a single LLM without our training recipe fails to improve accuracy. Moreover, Ex7 underperforms compared to training diffusion alone without hybrid loss (Ex2). The various token formulation designs used in our training recipe are explored in Table 1. The impact of denoising steps. In Figure 3, we explore the relationship between the manipulation performance and different denoising steps, comparing Ex1 (HybridVLA-dif), Ex2, and Ex5. The results suggest that integrating the diffusion process into the LLM and leveraging large-scale robotic pretraining allows for a reduction in inference denoising steps from 20 to 4 without causing performance degradation. Conversely, Ex5 (without large-scale pretraining) exhibits a significant decline in manipulation accuracy. To balance inference speed and accuracy, we set the diffusion denoising steps to 4. In Appendix B.2, we provide additional ablation studies on (1) confidence thresholds in the collaborative action ensemble and (2) KV cache influence on inference speed.
Franka single-arm robot AgileX dual-arm robot
Pick Unplug Pour Wipe Open drawer Mean. Pick Lift ball Place bottles Wipe Fold Mean.
Models and place charger water blackboard and place inside S.R. $\uparrow$ and place and place at rack blackboard shorts S.R. $\uparrow$
$\pi_{0}$ (2.6B) [8] 0.50 0.35 0.45 0.35 0.60 0.45 0.75 0.65 0.40 0.30 0.65 0.55
CogACT (7B) [52] 0.80 0.70 0.40 0.65 0.50 0.61 - - - - - -
HybridVLA-dif(7B) 0.85 0.95 0.75 0.85 0.60 0.80 0.80 0.75 0.60 0.45 0.70 0.66
Table 4: Comparison of our method and baselines in real-world scenarios. We train all methods in a single-task setting [110] and report the success rates. Success is determined by human evaluation based on whether the task is completed. Since CogAct lacks support for multi-view images, which are crucial for dual-arm tasks [8, 20], we conduct our dual-arm comparison solely with $\pi_{0}$.
4.3 Real-World Experiment
Self-collected Data. For single-arm tasks, we use a Franka Research 3 robot with a static front-view and a wrist-view camera. We perform 5 tasks: 1) Pick and place, 2) Unplug charger, 3) Open drawer and place inside, 4) Pour water, 5) Wipe blackboard. For each task, 100 demonstrations are collected via teleoperation using a SpaceMouse from various positions on the table. For dual-arm tasks, we use an AgileX dual-arm robot equipped with a static exterior view, a right-wrist view, and a left-wrist view camera. We conduct 5 coordinated dual-arm tasks: 1) Pick and place, 2) Lift ball and place, 3) place two bottles at rack, 4) Wipe blackboard, 5) Fold shorts. Similarly, 100 demonstrations are collected for each task using master-puppet teleoperation. Additional details are provided in Appendix A.2.
Training and Evaluation Details. For a fair comparison, we evaluate only the diffusion-generated actions of our method (HybridVLA-dif) against diffusion-based VLA methods, $\pi_{0}$[8] and CogAct[52]. The implementation details remain consistent with our simulation experiments, except for incorporating two-view inputs for single-arm tasks and three-view inputs for dual-arm tasks. For evaluation, we use the checkpoint from the latest epoch to perform 20 rollouts across diverse tabletop positions.
Figure 3: The impact of denoising steps, where the x-axis and y-axis represent the denoising steps and manipulation success rate.
Quantitative Results. In Table 4, HybridVLA-dif achieves outstanding performance across single-arm real-world tasks. For Pick and place and Unplug charger, our method achieves success rates of 85% and 95%, respectively, demonstrating accurate object position prediction. For Pour Water, HybridVLA-dif outperforms the previous SOTA method by 30%, demonstrating its ability to comprehend object relationships and predict precise rotations. The superior performance in Wipe Blackboard and Open Drawer and Place Inside further underscores HybridVLA-dif’s stability in long-horizon tasks. For dual-arm tasks, we extend the action dimension of diffusion tokens to 14-DOF, representing the 7-DOF end-effector poses for both the right and left arms. Our method consistently surpasses previous VLA method across 5 distinct tasks, highlighting HybridVLA’s ability to leverage VLMs’ reasoning capabilities for object affordance understanding and dual-arm coordination.
Qualitative Results. In Figure 4, we present visualizations of the manipulation process performed by our method. HybridVLA-dif accurately predicts control actions across various task demands, including precise positioning and rotation, dual-arm coordination, and scene understanding. For instance, in the Wipe blackboard dual-arm task, our method enables the left arm to precisely grasp the eraser while the right arm holds the edge of the blackboard. It then coordinates the movement and rotation of both objects to establish a relative spatial relationship. Finally, the left arm controls the eraser to precisely remove the red circle. Additional failure case analyses are provided in Appendix D, and execution videos are available in the supplementary materials.
Figure 4: The qualitative results of HybridVLA-dif in real-world tasks. More visualizations are shown in Appendix C.
Task Pick and place(single arm) Lift ball and place(dual arm)
Scenario HybridVLA-dif Cogact HybridVLA-dif $\pi_{0}$
Original 0.85 0.80 0.75 0.65
Object 0.50(-41%) 0.45(-43%) 0.70(-7%) 0.60(-8%)
Background 0.60(-31%) 0.50(-37%) 0.60(-20%) 0.50(-23%)
Height 0.70(-17%) 0.50(-37%) 0.55(-20%) 0.45(-31%)
Lightning 0.70(-17%) 0.60(-25%) 0.65(-13%) 0.55(-15%)
Table 5: Generalization. “Object”, “Background”, “Height”, and “Lighting” denote unseen manipulated objects, backgrounds, spatial positions, and lighting conditions, respectively. The image above illustrates the four unseen test scenarios, with red boxes highlighting the key differences.
4.4 Generalization Experiment
Beyond achieving stable manipulation across diverse tasks, HybridVLA also exhibits strong generalization capabilities in real-world scenarios. Since CogAct and $\pi_{0}$ excel in single-arm and dual-arm tasks, respectively, we design four common generalization experiments, comparing our method with CogAct on the single-arm Pick and place task and with $\pi_{0}$ on the dual-arm Lift ball and place task. 1) Unseen manipulated objects. In this scenario, we replace the training manipulated objects with a series of unseen objects, e.g., replacing the red block with a charger. As shown in the “Object” row of Table 5, our method demonstrates the smallest accuracy drop. These results indicate that HybridVLA effectively incorporates diffusion into next-token prediction, leveraging VLM pretraining knowledge to reason about diverse object semantics based on observation and language context. 2) Unseen background. In this scenario, cluttered backgrounds are introduced during testing, such as adding unseen flowers around the manipulated object. As shown in the “Background” row of Table 5, HybridVLA exhibits minimal accuracy degradation. These results demonstrate that our collaborative training recipe effectively absorbs the strengths of both action generation methods, enhancing robustness to environmental variations. 3) Unseen Spatial position. Unlike position shifts within the same plane, we introduce height variations during testing, further challenging the model’s spatial comprehension. As shown in the “Height” row of Table 5, HybridVLA consistently achieves precise manipulation even when encountering objects in previously unseen spatial positions. These results highlight that HybridVLA not only excels in observational generalization but also exhibits strong trajectory generalization capabilities. 4) Unseen lighting conditions. Finally, we introduce variations in lighting conditions, a common challenge in real-world environments. As shown in the last row of Table 5, all methods maintain satisfactory performance. This can be attributed to large-scale pretraining on robotic datasets, which significantly enhances these models’ ability to generalize across diverse data distributions.
5 Conclusion and Limitation
In this paper, we introduce HybridVLA, a unified framework that equips a single LLM with both diffusion and autoregressive action prediction capabilities. Our proposed collaborative training recipe bridges the gap between these two action generation approaches, enabling mutual reinforcement and enhancing manipulation robustness. By effectively integrating the inherently continuous nature of actions from diffusion with the reasoning capabilities of autoregressive methods, HybridVLA achieves outstanding performance and strong generalization across both simulation and real-world tasks. One limitation of HybridVLA is that its inference time is constrained by the slower autoregressive generation, reducing control frequency, similar to previous autoregressive-based VLA methods [47, 10, 56]. However, our collaborative training enables mutual reinforcement between the two generation methods, allowing inference with only the diffusion process (HybridVLA-dif) and achieving a 9.4 Hz inference speed.
Appendix A. We begin by detailing the large-scale pretraining and self-collected real-world datasets.
Appendix B. Additional simulation experiments and ablation studies are presented.
Appendix C. We include further visualizations of both single-arm and dual-arm manipulation processes.
Appendix D. An analysis of failure cases encountered when using HybridVLA to control a robot.
Appendix A Additional Dataset Details
A.1 Large-scale Pretraining Dataset
Our pre-training dataset collection comprises 35 datasets, encompassing a total of 760k trajectories and 33m frames. Table 6 provides a comprehensive list of our pre-training datasets along with their respective sampling weights. The number of trajectories and the sampling weights can be automatically adjusted during dataset assembly. Following the prior data preprocessing approach [47], we reformulate the pre-training datasets to emphasize end-effector sequence control, ensuring alignment with the specific requirements of our model training. Due to inherent differences among datasets, only single 2D observations are used during pre-training. However, during fine-tuning, HybridVLA can accommodate both single- and multi-view observations depending on the task requirements. For instance, AgileX dual-arm robot tasks require three viewpoints—an ego view and two wrist camera views—to capture a comprehensive observation of the object while mitigating occlusions caused by the robot arm. HybridVLA processes multi-view images using a shared vision encode and then concatenates the visual feature along the token dimension. Notably, the difference in the number of images used during pre-training and fine-tuning does not impact manipulation performance in downstream tasks.
Training Dataset Mixture
Fractal [9] 9.1%
Kuka [43] 27.8%
Bridge[19, 95] 4.1%
Taco Play [81, 66] 2.1%
Jaco Play [17] 0.3%
Berkeley Cable Routing [62] 0.2%
Roboturk [65] 1.7%
Viola [120] 0.7%
Berkeley Autolab UR5 [13] 0.9%
Toto [117] 1.5%
Language Table [64] 3.1%
Stanford Hydra Dataset [6] 3.2%
Austin Buds Dataset [119] 0.2%
NYU Franka Play Dataset [16] 0.6%
Furniture Bench Dataset [29] 1.8%
UCSD Kitchen Dataset [108] $<$0.1%
Austin Sailor Dataset [68] 1.6%
Austin Sirius Dataset [57] 1.2%
DLR EDAN Shared Control [76] $<$0.1%
IAMLab CMU Pickup Insert [83] 0.7%
UTAustin Mutex [85] 1.6%
Berkeley Fanuc Manipulation [118] 0.6%
CMU Stretch [67] 0.1%
BC-Z [37] 5.4%
FMB Dataset [63] 5.0%
DobbE [84] 1.0%
DROID [46] 7.2%
Stanford Kuka Dataset [49] 0.1%
Stanford Robocook Dataset [86] 0.1%
Maniskill [26] 6.3%
Berkeley RPT [78] 0.1%
QUT Dexterous Manipulation [11] 0.1%
RoboSet [48] 1.8%
BridgeData V2 [95] 4.7%
RoboMind [101] 5.2%
Table 6: The dataset name and sampling weight used in our mixed large-scale pretraining dataset.
Figure 5: Real-World Assets and Experimental Settings. We provide visualizations of the assets used and the experimental settings for single-arm FR3 robot tasks and dual-arm AgileX robot tasks, respectively.
Close Close Toilet Sweep Close Phone Umbrella Frame Wine at Water Mean.
Models box laptop lid seat down to dustpan fridge on base out off hanger rack plants S.R. $\uparrow$
HybridVLA-ar(7B) 0.85 0.70 0.90 0.85 0.95 0.30 0.25 0.40 0.45 0.50 0.62
HybridVLA-dif(7B) 0.85 0.75 1.0 0.80 0.95 0.50 0.50 0.30 0.70 0.25 0.66
HybridVLA(7B) 0.85 0.95 1.0 0.90 1.0 0.50 0.50 0.70 0.50 0.50 0.74
Table 7: Detailed Simulation Experiments. We validate that different action types within our proposed framework exhibit varying performance across tasks. All models undergo joint training using our proposed collaborative training recipe; however, HybridVLA-ar and HybridVLA-dif rely exclusively on autoregressive-based and diffusion-based action generation during inference, respectively. Underlining indicates the highest score between HybridVLA-ar and HybridVLA-dif.
Threshold 0.90 0.92 0.94 0.96 0.98
Success rate 0.66 0.64 0.70 0.74 0.69
Table 8: Ablation Study. We explore the impact of different confidence thresholds on the performance of ensemble actions.
A.2 Self-collected Real-world Dataset
The experimental assets and environments for the single-arm and dual-arm setups are shown in Figure 5 (a) and (b), respectively. For the single-arm setup, a 3D-printed UMI gripper [15] is attached to the Franka robot and is used across all baselines. We utilize RealSense 435 and RealSense 515 cameras to capture both wrist and front views. For the dual-arm setup, two Orbbec DABAI cameras are used to capture the left and right wrist views, while a RealSense 515 is mounted overhead to capture a static third-person view. We provide a detailed explanation of the real-world tasks and their success conditions. We begin by describing the single-arm tasks:
1. Pick and place. This task requires the robot to pick up a specifically colored block based on a language description and place it in a specifically colored bowl.
2. Unplug charger. The robot needs to grasp the charger at an optimal position and rotation, and then lift it to a certain height without slipping.
3. Pour water. The robot needs to first pick the bottle, then rotate it to a position slightly above the cup, and tilt it to perform the pouring action. The task is deemed successful only if the bottle opening is correctly aligned with the cup.
4. Wipe blackboard. The robot needs to first grasp an eraser and then use it to remove the red markings from a blackboard placed on the tabletop. The red markings are drawn on an unfixed region, and the task is considered successful only if they are completely erased.
5. Open drawer and place inside. The robot needs to open the top drawer, pick up the required objects based on the language description, place them in the opened drawer, and then close it. This task consists of four sequential sub-tasks: open drawer, pick object, place object, and close drawer. The task is considered complete once all sub-tasks have been successfully executed.
We then describe the details of dual-arm tasks:
1. Pick and place. The robot must use both its left and right arms to pick up two objects based on the language description and place them in the container.
2. Lift ball and place. Both the left and right arms must simultaneously make contact with the ball, which is secured between the two grippers. The arms coordinate their movements to transport the ball to the container while ensuring it does not slip. This task highly tests the model’s dual-arm coordination capabilities.
3. Place bottles at rack. The left and right robot arms need to grasp the bottles placed on their respective sides and rotate them to position them parallel to the rack.
4. Wipe blackboard. Unlike the single-arm setting, the dual-arm setting requires one arm to hold the whiteboard while the other picks up the eraser and wipes off the red marker.
5. Fold shorts: This task requires folding a pair of shorts, involving two sequential steps. First, one pant leg is folded over the other to align them. Then, the pants are folded in half from top to bottom. Throughout the process, both arms must coordinate their movements. For example, in the first step, the left arm holds the bottom of the pant leg while the right arm grips the upper part, working together to complete the folding.
Figure 6: Single-arm Execution Visualization. We visualize key frames of the agent’s execution process from the front perspective.
Appendix B Additional Quantitative Results
B.1 Additional Simulation Experiments
In Table 7, we validate the first observed phenomenon mentioned in Section 3.3: different action types within our proposed framework exhibit varying performance across tasks. Meanwhile, we categorize our method into three modes: HybridVLA (7B), HybridVLA-ar (7B), and HybridVLA-dif (7B). All modes undergo joint training using our proposed collaborative training recipe; however, HybridVLA-ar and HybridVLA-dif rely exclusively on autoregressive-based and diffusion-based action generation during inference, respectively. The experiments are conducted in the RLBench simulator across 10 tasks, and evaluated based on success rate. Comparing HybridVLA-ar and HybridVLA-dif, HybridVLA-ar outperforms in 4 out of 10 tasks, while HybridVLA-dif leads in the remaining 6 tasks. These results validate our findings that, within the HybridVLA framework, diffusion-based predictions excel in precise manipulation tasks, such as Phone on base, Toilet seat down, and Close laptop lid, whereas autoregressive predictions perform better in tasks requiring scene-level semantic reasoning, such as Sweep to dustpan, Water plants, and Frame off hanger. Therefore, while collaborative training allows diffusion-based and autoregressive-based action generation to reinforce each other, assembling both methods results in more robust actions.
Figure 7: Dual-arm Execution Visualization. We visualize key frames of the agent’s execution process from a static exterior view.
B.2 Additional Ablation Study
The impact of confidence threshold in collaborative action ensemble. The proposed collaborative ensemble strategy determines whether to use the action predicted by diffusion alone or the averaged output of both diffusion and autoregressive methods, guided by a mean confidence threshold derived from the autoregressive action token. In this experiment, we investigate the optimal confidence threshold required to ensure the accuracy of autoregressive actions and enhance the overall precision of the ensemble-generated action. Specifically, as shown in Tab. 8, we vary the threshold from 0.90 to 0.98. We find that when the confidence threshold drops below 0.94, autoregressive predictions become unreliable, leading to a slight degradation in the performance of the ensemble action. Conversely, when the threshold reaches 0.98, the number of valid autoregressive actions becomes too limited, causing the performance of the ensemble action to closely match that of the diffusion-predicted action. Empirically, we conclude that setting the threshold to 0.96 ensures a stable action ensemble.
The impact of KV cache in inference speed. As described in Section 3.3, we adopt the KV cache to eliminate redundant computations and enhance inference speed. In this experiment, we examine the extent to which this mechanism accelerates inference. With the KV cache enabled (Table 2 of the main paper), HybridVLA-dif achieves an average success rate of 66% across 10 simulation tasks with an inference speed of 9.4 Hz. Removing it results in a similar average success rate but reduces the inference speed to 5.0 Hz. While the KV cache has typically been used in previous autoregressive VLA methods [47, 56], we are the first to integrate it into an LLM’s diffusion-based action generation.
Appendix C Additional Visualizations
Figure 6 and Figure 7 illustrate keyframes of single-arm and dual-arm real-world execution processes. Notably, our Franka Research 3 (FR3) operates with controller version 5.6.0, libfranka version 0.13.3, Franka ROS version 0.10.0, and Ubuntu 20.04 with ROS Noetic. Under these software settings, the FR3 remains in green light execution mode with the FCI switch set to ‘on’.
These tasks demonstrate HybridVLA’s capability in accurately predicting position and rotation, as well as determining the precise timing for changing the gripper’s open state. Additionally, the dual-arm tasks highlight HybridVLA’s ability to coordinate both robotic arms, enabling it to complete tasks beyond the capability of a single arm, such as transporting a ball to a container. Notably, the single-arm task ‘open drawer and place’ and the dual-arm tasks ‘wipe whiteboard’ and ‘fold shorts’ are long-horizon tasks that involve at least three multi-step actions. These results further confirm that HybridVLA can reliably predict sequential actions, demonstrating the capability to complete long-horizon tasks.
Appendix D Failure Case Analysis.
Through extensive real-world experiments, we identify three primary failure categories that impact the performance of HybridVLA. The first category, rotational prediction deviations, is particularly evident in tasks requiring precise rotation control, such as Pour water and Place bottle at rack. These failures include accumulated errors in multi-step rotational movements and incorrect rotation angles when interacting with target objects. The second category pertains to pose predictions that exceed the robot’s degree of freedom limits. The model sometimes predicts poses beyond the mechanical constraints of the Fr3 arm or AgileX dual-arm robot, generates target positions that fall outside the workspace boundaries, or produces kinematically infeasible configurations during complex transitions. The third category involves failures in dual-arm coordination, where both arms must collaborate to complete a task. Since the model predicts each arm’s actions based on the current object state, any interaction by one arm can alter the object’s state, potentially invalidating the previously predicted action of the other arm.